Instructions say: "Find an equation in standard form for the line described." Here's the question: "Through P (-2, 1) and parallel to the line containing (1, 4), (2, 3). I have no idea. Could you please help?
Answer by stanbon(75887)

Instructions say: "Find an equation in standard form for the line described." Here's the question: "Through P (-2, 1) and parallel to the line containing
(1, 4), (2, 3).
--------
The slope thru (1,4) and (2,3) is (3-4)/(2-1) = -1
----
Any line parallel to that line must have a slope of -1
----
Equation of the line you want:
x = -2, y = 1, m = -1
y = mx + b; solve b
----
1 = -1*-2 + b
b = -1
----
Equation:
y = -x-1
==================
Cheers,
Stan H.
===========